# taskpriors: Enhancing Model Evaluation by Considering the Entire Space of Downstream Tasks

### v2

@@ Abstract
The grand goal of AI research, and particularly Self Supervised Learning (SSL), is to produce systems that can successfully solve any possible task. In contrast, current evaluation methods available to AI researchers typically rely on a fixed collection of hand-picked downstream benchmarks. Hence, a large amount of effort is put into designing and searching for large collection of evaluation tasks that can serve as a proxy of our grand goal. We argue that such a rigid evaluation protocol creates a silent bottleneck in AI research. To remedy that, we define a probabilistic space of downstream tasks obtained by adopting a distribution of tasks and by defining Task Priors [^1] https://github.com/niketp03/taskpriors, see Appendix C for a simple example.. Under this view, one can evaluate a model’s performance over the set of all possible downstream tasks. Our framework is the first to provide answers to key questions such as (i) what is the average performance of my model over all possible downstream tasks weighted by the probability to encounter each task? or (ii) what is the variance of my model’s performance across all downstream tasks under the defined Task Priors? Beyond establishing a new standard for evaluation, we believe that Task Priors will accelerate the pace of research in SSL–where downstream task evaluation is the sole qualitative signal that researchers have access to.
@@ 1 Introduction
Pretrained backbone models today are released as a single checkpoint, yet in practice they can power millions of distinct downstream tasks: from simple classification and retrieval services to large-scale recommendation, autonomous perception, and more. On HuggingFace alone, top models such as mobilenetv3, clip-vit, and bert can get over 100 million downloads per month [27]. As the number of users and the diversity of applications grows, the space of possible downstream tasks users want their models to perform well on tends toward infinity.
Yet our standard evaluation protocols remain tethered to a small, fixed suite of hand-picked benchmarks—often half a dozen or so datasets (e.g., ImageNet, COCO, GLUE, SuperGLUE, WMT) [12, 19, 16, 25]. Each new benchmark can take months of expert labeling and tens or even hundreds of thousands of dollars to assemble. Even large-scale benchmark suites that aggregate many tasks — such as the Massive Text Embedding Benchmark (MTEB), which spans 56 different evaluation datasets — still represent only a narrow slice of the possible task space [18]. Once built, a static benchmark can only ever probe a tiny corner of the real-world tasks for which a model might be deployed. This disconnect creates a structural bottleneck between the handful of evaluation suites we all agree to evaluate our models on and the effectively infinite variety of tasks practitioners use our models for.
One way to break this bottleneck is simply to keep spending more time and money on ever-larger benchmarks, but that approach quickly becomes unsustainable. Instead, we propose a surrogate evaluation framework that treats downstream tasks as samples from a well-defined probabilistic space. By adopting a “Task Prior”, a distribution over all possible targets informed by a pretrained feature kernel, we can compute expectations and variances of a model’s downstream performance in closed form, without training new classifiers or designing new benchmarks. We show that this approach provides principled measures of average performance, robustness, and worst-case error across the entire landscape of potential tasks.
@@
  • Task Priors. We prove that performance on linear probes is equivalent to maximizing the alignment of a kernel with the induced label graph, unifying supervised and self-supervised evaluation. We then formalize the space of downstream tasks as a Gibbs-style distribution over all possible label graphs, making the notion of “any task a user might care about” mathematically explicit.
  • Closed-form Metrics. From the Task Prior we derive $O(n^{2})$ formulas for the expected downstream error and its variance, obviating benchmark curation or probe training.
  • Efficient Task Sampling. We introduce an $O(n)$ prefix-sampling algorithm that draws realistic classification tasks from the prior, enabling cheap evaluation on many tasks.
  • Empirical Validation. We find that across a range of backbones, our closed form metrics on kernel align with mean and variance of accuracy of linear probes on a subset of ImageNet. We then introduce a training objective that maximizes this expected task performance while simultaneously minimizing its variance across tasks.
2 Background
Many works aim to capture the performance of representation models primarily by intrinsic quantities about the model’s features. For instance, RankMe measures the effective rank of the feature matrix and shows a empirical correlation with average linear-probe accuracy across several tasks [7]. LiDAR argues that a method built on Linear Discriminant Analysis serves as a proxy for downstream performance [23]. More recent works take a broader view, demonstrating that k-NN, few-shot fine-tuning, and clustering evaluations may all disagree in systematic ways [17]. Collectively, these studies show that properties intrinsic to the representation can forecast downstream success, but they still reduce performance to one or two scalar summaries.
A complementary line of work attacks the evaluation bottleneck by increasing the number of test tasks. In NLP, suites such as MTEB (56 embedding datasets) [18] and HELM (42 scenarios, seven axes of measurement) [15] provide broad coverage of downstream tasks. The same trend is exists in vision, with works such as VideoEval packaging twenty diverse video understanding datasets together [13]. While these mega-benchmarks can be quite helpful for practitioners, they remain finite and expensive to create. Worse, even a hundred benchmarks sample only a vanishingly small corner of the large task space practitioners can care about.
Our Task Priors framework can be viewed as the missing bridge between these two threads. Like intrinsic metrics, it avoids needing a hand curated set of downstream targets, but like conglomerate benchmarks, it explicitly reasons about many tasks. Our framework echoes several well-known results from the classical theory of kernels. Notably, the trace term in Theorem 3.3 parallels the Hilbert–Schmidt Independence Criterion (HSIC) of Eq. (4) in [8]. Likewise, the term we get by taking the trace of ${\mathbf{K}}{\mathbf{G}}$ is precisely the same as “kernel alignment” studied in the context of generalization [5], obtained by flattening each matrix and taking their inner product, as $\langle{\mathbf{K}},{\mathbf{G}}\rangle=\operatorname{Tr}({\mathbf{K}}{\mathbf{G}})$.
There are some other related works that attack similar problems. In the computer vision space, VTAB [30] takes a similar distributional view of tasks, but does not precisely characterize the distribution of tasks. Similar to our derivations, [14] proposes a loss function based on the HSIC, which is an interesting avenue for future research.
3 Task Priors
@@ 3.1 An Equivalence Between Losses
Traditional downstream evaluation asks, “How well can a linear head map features back to a specific label vector $\mathbf{Y}$?”. This works when a single, human-defined labeling is the goal, but it scales poorly once we care about many downstream tasks. By recasting supervised objectives in the language of pairwise relations by looking at the sample-sample adjacency matrix $\mathbf{G}$ we leave the “label space” and move into a kernel space where tasks become graphs and models can be represented in terms of the kernel matrix. This shift from comparing features to labels, to comparing features among themselves, allows for us to derive a natural equivalence between supervised and self-supervised objectives, such as those found in contrastive learning [24, 1, 9, 2, 3, 4, 10, 29, 26]. We will leverage this connection to define our Task Prior. To better explain our findings, we first need to introduce the formal definitions of relative and absolute objectives.
@@
In many practical settings, rather than focusing solely on self-supervised loss formulations, we are given a pretrained representation $f_{\theta}(\mathbf{X})$ and wish to assess how well a simple linear head can recover downstream tasks. By solving the unconstrained mean squared error objective for the linear probe, we find that the optimal loss depends only on the kernel matrix ${\mathbf{M}}=f_{\theta}(\mathbf{X})^{\top}f_{\theta}(\mathbf{X})$ and the label adjacency graph ${\mathbf{G}}={\mathbf{Y}}^{\top}{\mathbf{Y}}$:
@@
Our goal will be to minimize that loss with respect to ${\mathbf{W}},{\mathbf{b}}$ and to perform some algebraic manipulations to demonstrate how that absolute objective–comparing the prediction of sample $n$ to the label of sample $n$–turns into a relative objective–comparing pairwise predictions to pairwise labels. The following derivations do not rely on any assumptions about the input ${\mathbf{X}}$, the network $f_{{\mathbf{\theta}}}$, or the targets ${\mathbf{Y}}$, other then that the representations $f_{{\mathbf{\theta}}}({\mathbf{X}})$ are of full rank, which is generally true. We will also denote the Singular Value Decomposition of $f_{\theta}({\mathbf{X}})$ as ${\mathbf{U}}{\mathbf{\Sigma}}{\mathbf{V}}^{\top}$, and the diagonal of ${\mathbf{\Sigma}}$ as $\sigma_{1},\dots,\sigma_{K}$.
@@
(Proof in section A.1.)
This theorem shows that the absolute objective from 3.1 is equivalent to the associated relative objective in 3.2. A crucial benefit of eq. 2 is that it only requires ${\mathbf{G}}$ which could be built from ${\mathbf{Y}}$–or without labels at all. For example, in a classification setting, definition 3.1 requires the categorical label for each sample, while definition 3.2 only requires to know which samples are from the same class but not what that class is. It is clear that building ${\mathbf{G}}$ directly instead of ${\mathbf{Y}}$ is easier and cheaper to get as it reduces the amount of expert knowledge needed.
In the relative objective, one no longer tries to predict ${\mathbf{y}}_{n}$ from ${\mathbf{x}}_{n}$, but instead tries to make the pairwise comparison of predictions ${\mathbf{V}}{\mathbf{V}}^{\top}$ aligned with the pairwise comparison of labels ${\mathbf{Y}}^{\top}{\mathbf{Y}}$. In particular, we clearly see that the left singular vectors of $f_{{\mathbf{\theta}}}({\mathbf{X}})$ no longer contribute to the loss–as expected since the optimal ${\mathbf{W}}$ automatically maps them to the left singular vectors of ${\mathbf{Y}}$. Hence, we are left with the right singular vectors ${\mathbf{V}}$ of $f_{{\mathbf{\theta}}}({\mathbf{X}})$ and its singular values that appear within ${\mathbf{D}}$. Notice that the trace on the right hand side of (2) is not trivial, as $\mathbf{V}$ denotes the truncated SVD, and has orthogonal columns but does not necessarily have orthogonal rows. This is especially true when the number of data points exceeds the feature dimension. So, since we can regard ${\mathbf{Y}}^{\top}{\mathbf{Y}}={\mathbf{G}}$ and ${\mathbf{V}}{\mathbf{V}}^{\top}$ is a valid kernel we can write as ${\mathbf{K}}$, then we can write our trace as $-\frac{1}{N}\operatorname{Tr}({\mathbf{G}}{\mathbf{K}})$.
Key Takeaway: Objectives that compare samples amongst themselves in a graph structure emerge from considering supervised objectives that compare sample predictions to labels.
@@ 3.2 A Distribution Over Target Labels
Building on the kernel view from Theorem 3.3, we now introduce a prior distribution over label graphs ${\mathbf{G}}$ that reflects the likelihood of different downstream tasks. Instead of manually selecting or training probes on individual benchmarks, we define a measure over all possible $\mathbf{G}$ matrices, weighted by alignment with a pretrained feature kernel $\mathbf{K}$. This Task Prior allows us to compute expectations and variances of kernel alignment scores in closed form, or to efficiently sample realistic downstream targets without training additional classifiers. To achieve this, a practical solution is to define a distribution over possible labelings which assigns a higher probability to more realistic scenarios that might be encountered in real-world applications. This will allow us to look into the average performance of our model, but also have the ability to look into other statistical properties.
(a) TSNE of Imagenette [11] with standard labels.
(b) Labels Generated by the Task Prior. [^3] Interested readers can experiment with different class counts and temperature settings in this Colab notebook: https://colab.research.google.com/drive/1qNOgoNSH87AcdODug-yop7Q0MuT8w1r7
Figure 1: Comparison of the naive way to evaluate a model, only on the specific choice of labels provided with the dataset (Left) with the probabilistic view of targets generated by sampling from the Task Prior, giving us a distribution we can evaluate on (Right).
Suppose that we have a kernel function that measures similarity between data points $k:\mathbb{R}^{D}\times\mathbb{R}^{D}\to\mathbb{R}$. Let $\mathbf{K}$ then be a kernel matrix corresponding to ${\mathbf{X}}$. For the rest of the paper we can assume this is the centered kernel matrix corresponding to the features. Now if we have a graph adjacency matrix ${\mathbf{G}}$, we can read off the elements of the product,
$[\mathbf{G}\mathbf{K}]_{ij}=\sum_{k=1}^{N}\mathbf{G}_{ik}\ k({\mathbf{x}}_{j},{\mathbf{x}}_{k}).$
In particular we can see that the diagonal elements of this matrix is given as follows, where we use ${\mathbf{x}}_{j}\sim{\mathbf{x}}_{i}$ to mean that ${\mathbf{x}}_{j}$ and ${\mathbf{x}}_{i}$ are connected in graph $\mathbf{G}$,
$[\mathbf{G}\mathbf{K}]_{ii}=\sum_{k=1}^{N}\mathbf{G}_{ik}\ k({\mathbf{x}}_{i},{\mathbf{x}}_{k})=\sum_{{\mathbf{x}}_{k}\sim{\mathbf{x}}_{i}}k({\mathbf{x}}_{k},{\mathbf{x}}_{i}).$
Summing over $i$ gives the trace,
@@
which acts as a global compatibility score between the geometry in the label graph $\mathbf{G}$ and the kernel $\mathbf{K}$. We treat the negative trace, $\mathcal{E}(\mathbf{G}):=-\operatorname{Tr}(\mathbf{G}\mathbf{K}),$ as the ”energy“ of a labeling: graphs that connect feature-similar points (high trace of $\mathbf{G}\mathbf{K}$) have lower energy and are therefore more likely. This leads to the naturally to the Gibbs measure
$\mu(\mathbf{G})\;\propto\;e^{-\mathcal{E}(\mathbf{G})/T}\;=\;e^{\operatorname{Tr}(\mathbf{G}\mathbf{K})/T},$
with temperature $T>0$. As we increase the temperature, this distribution tends towards the uniform distribution on all graphs, and we get more interesting behavior in the low temperature regime. The properties of this distribution enable direction computation of the expectation and higher moments, and enable efficient sampling algorithms which we develop in the sequel.
@@
(Task Prior Distribution) Given a kernel matrix $\mathbf{K}$ on $n$ data points, and a temperature $T>0$, we will define the following Gibbs measure on the space of all graphs, $\mathbf{\mathbf{missing}}G$ as:
@@
Although computing exactly the probability of observing a single graph can be quite challenging, as computing the partition function would require $2^{N^{2}}$ computations, the specific structure of this probability measure admits a neat factorization on a per-edge level.
@@
Key Takeaway: Modeling label graphs with a Gibbs prior aligned to a backbone’s feature kernel, turns the space of graphs on the data points into a tractable distribution, yielding closed form calculations of the mean and variance.

### v3

@@ Abstract
A long-standing research problem in Artificial Intelligence (AI) is to produce systems that can successfully solve any possible task. A key requirement in addressing progress in that direction is a near-infinite suite of tasks for benchmarking AI solutions. In contrast, current evaluation methods available to AI researchers in representation learning typically rely on a fixed collection of hand-picked downstream benchmarks. Hence, a large amount of effort is put into designing and searching for large collection of evaluation tasks that can serve as a proxy of our grand goal. We argue that such a rigid evaluation protocol creates a structural bottleneck in AI research. To remedy that, we define a probability distribution over downstream tasks – Task Priors. Under this view, one can evaluate a model’s performance over the set of all possible downstream tasks. Our framework is the first to provide answers to key questions such as (i) what is the average performance of my model over all possible downstream tasks weighted by the probability to encounter each task? or (ii) what is the variance of my model’s performance across all downstream tasks under the defined Task Priors? Beyond establishing a new standard for evaluation, we believe that Task Priors will accelerate the pace of research in representation learning – where downstream task evaluation is generally the sole signal that researchers have access to.
Figure 1: Comparison of the naive way to evaluate a model, only on the specific choice of labels provided with the Imagenette (Howard, 2019) dataset (Left) with the probabilistic view of targets sampled from the Task Prior, giving us a distribution we can evaluate on (Right).
@@ 1 Introduction
Pretrained backbone models today are released as a single checkpoint, yet in practice they can power millions of distinct downstream tasks: from simple classification and retrieval services to large-scale recommendation, autonomous perception, and more. On HuggingFace alone, top models such as mobilenetv3, clip-vit, and bert receive over 100 million downloads per month (Wightman, 2019). As the number of users and the diversity of applications grows, the space of possible downstream tasks that users want their models to perform well on tends toward infinity.
Yet, our standard evaluation protocols remain tethered to a small, fixed suite of hand-picked benchmarks, often around less than ten datasets (e.g., ImageNet, COCO, GLUE, SuperGLUE, WMT) (Lhoest et al., 2021; Russakovsky et al., 2015; Lin et al., 2014; Wang et al., 2018). Each new benchmark can take months of expert labeling and tens or even hundreds of thousands of dollars to assemble. Even large-scale benchmark suites that aggregate many tasks, such as the Massive Text Embedding Benchmark (MTEB), which spans 56 different evaluation datasets, or the similar Massive Image Embedding Benchmark (MIEB), still represent only a narrow slice of the possible space of downstream tasks (Muennighoff et al., 2023; Xiao et al., 2025). Once built, a static benchmark can only ever probe a tiny corner of the real-world tasks for which a model might be deployed. This disconnect creates a structural bottleneck between the handful of evaluation suites we all agree to evaluate our models on and the effectively infinite variety of tasks practitioners use our models for.
One way to break this bottleneck is simply to keep spending more time and money on ever-larger benchmarks, but that approach quickly becomes unsustainable. Instead, we propose a proxy evaluation framework that treats downstream tasks as samples from a well-defined probability space. By adopting a “Task Prior”, a distribution over all possible targets informed by a feature kernel, we can compute expectations and variances of a model’s downstream performance in closed form, without training new classifiers or designing new benchmarks. We show that this approach provides principled measures of average performance, robustness, and worst-case error across the entire landscape of potential tasks.
@@
  • Task Priors. We motivate the definition of the Task Prior by first proving that performance on linear probes is equivalent to maximizing the alignment of a kernel with the induced label graph, unifying supervised and self-supervised evaluation. We then formalize the space of downstream tasks as a Gibbs distribution over all possible label graphs.
  • Closed-form Statistics. From the Task Prior we derive $O(n^{2})$ formulas for the expected downstream error and its variance, bypassing benchmark curation or probe training.
  • Efficient Sampling from the Task Prior. We introduce an $O(n)$ prefix-sampling algorithm that draws realistic classification tasks from the prior, enabling cheap evaluation on many tasks.
  • Empirical Validation. We find that across a range of backbones, our closed form metrics on kernel align with mean and variance of accuracy of linear probes on a subset of ImageNet. We then show that the performance implied by our Task Prior correlates with performance on a hand curated set of downstream tasks.
2 Task Priors: Model Evaluation on Infinitely Many Tasks
This section bridges supervised “absolute” objectives and pairwise “relative” objectives and uses that link to define a Task Prior tied to our pretrained feature kernels. Concretely, we prove the loss equivalence (2.1), introduce a Gibbs prior over label graphs (2.2), connect it to SSL formulations (2.3), and present a fast sampler for evaluation of probes (2.4).
@@ 2.1 An Equivalence Between Losses
Traditional downstream evaluation relies on absolute objectives, and asks, “How well can a linear head map features back to a specific label vector $\mathbf{Y}$?”. This works when a single, human-defined labeling is the goal, but it scales poorly once we care about many downstream tasks. By recasting supervised objectives in the language of pairwise relations by looking at the sample-sample adjacency matrix $\mathbf{G}$, we can shift to look at the more general setting of relative objectives. In this section, we will relate these two types of objectives and derive a natural equivalence between supervised and self-supervised objectives. In the sequel, we will leverage this connection to define our Task Prior. To better explain our findings, we first need to introduce the formal definitions of relative and absolute objectives.
@@
In many practical settings, rather than focusing solely on self‑supervised loss formulations, we are given a pretrained representation $f_{\theta}(\mathbf{X})$ and wish to assess how well a simple linear head can recover downstream tasks. By solving the unconstrained mean squared error objective for the linear probe, we find that the optimal loss depends only on the kernel matrix ${\mathbf{M}}=f_{\theta}(\mathbf{X})^{\top}f_{\theta}(\mathbf{X})$ and the label adjacency graph ${\mathbf{G}}={\mathbf{Y}}^{\top}{\mathbf{Y}}$:
@@
Our goal will be to minimize that loss with respect to ${\mathbf{W}},{\mathbf{b}}$ and to perform some algebraic manipulations to demonstrate how that absolute objective–comparing the prediction of sample $n$ to the label of sample nn–turns into a relative objective–comparing pairwise predictions to pairwise labels. The following derivations do not rely on any assumptions about the input $\mathbf{X}$}, the network $f_{{\mathbf{\theta}}}$, or the targets $\mathbf{Y}$}, other then that the representations $f_{{\mathbf{\theta}}}({\mathbf{X}})$ are of full rank, which is generally true. We will also denote the Singular Value Decomposition of $f_{\theta}({\mathbf{X}})$ as ${\mathbf{U}}{\mathbf{\Sigma}}{\mathbf{V}}^{\top}$.
@@
(Proof in Appendix A.1.)
This theorem shows that the absolute objective from 2.1 is equivalent to the associated relative objective in 2.2. A crucial benefit of eq. 2 is that it only requires $\mathbf{G}$}, which could be built from $\mathbf{Y}$}, or without labels at all. For example, in a classification setting, definition 2.1 requires the categorical label for each sample, while definition 2.2 only requires to know which samples are from the same class but not what that class is.
In the relative objective, one no longer tries to predict ${\mathbf{y}}_{n}$ from ${\mathbf{x}}_{n}$, but instead tries to make the pairwise comparison of predictions ${\mathbf{V}}{\mathbf{V}}^{\top}$ aligned with the pairwise comparison of labels ${\mathbf{Y}}^{\top}{\mathbf{Y}}$. In particular, we clearly see that the left singular vectors of $f_{{\mathbf{\theta}}}({\mathbf{X}})$ no longer contribute to the loss, as expected since the optimal $\mathbf{W}$} automatically maps them to the left singular vectors of $\mathbf{Y}$}. Hence, we are left with the right singular vectors $\mathbf{V}$} of $f_{{\mathbf{\theta}}}({\mathbf{X}})$. Notice that the trace on the right hand side of equation 2 is not trivial, as $\mathbf{V}$ denotes the truncated SVD, and has orthogonal columns but does not necessarily have orthogonal rows. This is true in the typical case when the number of data points exceeds the feature dimension. So, since we can regard ${\mathbf{Y}}^{\top}{\mathbf{Y}}={\mathbf{G}}$ and ${\mathbf{V}}{\mathbf{V}}^{\top}$ is a valid kernel we can write as $\mathbf{K}$}, then we can write our trace as $\frac{1}{N}\operatorname{Tr}({\mathbf{G}}{\mathbf{K}})$.
@@ 2.2 A Distribution Over Target Labels
Building on Theorem 2.3, we can now introduce a prior distribution over label graphs $\mathbf{G}$} that reflects the likelihood of different downstream tasks. To achieve this, we define a measure over all possible $\mathbf{G}$ matrices, weighted by alignment with a pretrained feature kernel $\mathbf{K}$. This will assign a higher probability to more realistic scenarios that might be encountered in real-world applications. This Task Prior allows us to compute expectations and variances of kernel alignment scores in closed form, and to efficiently sample realistic downstream targets without training additional classifiers.
To begin, suppose that we have a kernel function that measures similarity between data points $k:\mathbb{R}^{D}\times\mathbb{R}^{D}\to\mathbb{R}$. Let $\mathbf{K}$ then be a kernel matrix corresponding to $\mathbf{X}$}. For the rest of the paper we can assume this is the centered kernel matrix corresponding to the features. Now if we have a graph adjacency matrix $\mathbf{G}$}, we can read off the elements of the product, $\mathbf{G}\mathbf{K}]_{ij}=\sum_{k=1}^{N}\mathbf{G}_{ik}\ k({\mathbf{x}}_{j},{\mathbf{x}}_{k}).$ In particular we can see that the diagonal elements of this matrix is given as follows, where we use ${\mathbf{x}}_{j}\sim{\mathbf{x}}_{i}$ to mean that ${\mathbf{x}}_{j}$ and ${\mathbf{x}}_{i}$ are connected in graph $\mathbf{G}$, $\mathbf{G}\mathbf{K}]_{ii}=\sum_{k=1}^{N}\mathbf{G}_{ik}\ k({\mathbf{x}}_{i},{\mathbf{x}}_{k})=\sum_{{\mathbf{x}}_{k}\sim{\mathbf{x}}_{i}}k({\mathbf{x}}_{k},{\mathbf{x}}_{i}).$ Summing over $i$ gives the trace,
@@
which acts as a global compatibility score between the geometry in the label graph $\mathbf{G}$ and the kernel $\mathbf{K}$. We treat the negative trace, $\mathcal{E}(\mathbf{G}):=-\operatorname{Tr}(\mathbf{G}\mathbf{K}),$ as the ”energy“ of a labeling: graphs that connect feature-similar points (high trace of $\mathbf{G}\mathbf{K}$) have lower energy and are therefore more likely. This leads to the naturally to the Gibbs measure, $\mu(\mathbf{G})\;\propto\;e^{-\mathcal{E}(\mathbf{G})/T}\;=\;e^{\operatorname{Tr}(\mathbf{G}\mathbf{K})/T},$ with temperature $T>0$. As we increase the temperature, this distribution tends towards the uniform distribution on all graphs, and we expect to see more interesting behavior in the low temperature regime. The properties of this distribution enable direction computation of the expectation and higher moments, and enable efficient sampling algorithms which we develop in the sequel.
Figure 2: Here we show an example of using Task Priors to evaluate the DinoV2 (Oquab et al., 2023) family of models, with respect the the Task Prior kernel, dinov2-giant. We show the expectation and variance of $\operatorname{Tr}({\mathbf{K}}{\mathbf{G}})$ (Left), as well as the distribution of performance of linear probes on labelings sampled from the Task Prior (Right).
@@
(Task Prior Distribution) Given a kernel matrix $\mathbf{K}$ on $n$ data points, and a temperature $T>0$, we will define the following Gibbs measure on the space of all graphs, $\mathbf{\mathbf{}}G$ as:
@@
Although computing exactly the probability of observing a single graph can be quite challenging, as computing the partition function would require $2^{N^{2}}$ computations, the specific structure of this probability measure admits a convenient factorization on a per-edge level.